The sum of two complex numbers is -3 + 5i. If one of these complex numbers is 1-8i, what is the other number?

Answer by Alan3354(69443)

The sum of two complex numbers is -3 + 5i. If one of these complex numbers is 1-8i, what is the other number?
----------
a + bi is the other number.
--------
The real and imaginary parts are combined separately.
---
1 - 8i + a + bi = -3 + 5i
1 + a = -3
a = -4
-----
-8 + b = 5
b = 13
===========
--> -4 + 13i